Find the length of a rectangular lot with a perimeter of 72 m if the length is 6 m more than the width.
Answer by algebrahouse.com(1659)

w = width
w + 6 = length {length is 6 more than width}


Perimeter of a rectangle = 2(width) + 2(length)

2w + 2(w + 6) = 72 {substituted width and length into perimeter formula}
2w + 2w + 12 = 72 {used distributive property}
4w + 12 = 72 {combined like terms}
4w = 60 {subtracted 12 from each side}
w = 15 {divided each side by 4}
w + 6 = 21 {substituted 15, in for w, into w + 6}

width = 15 m
length = 21 m

For more help from me, visit: www.algebrahouse.com